# One Model, Any Conjunctive Query: Graph Neural Networks for Answering Queries over Incomplete Knowledge Graphs

### v2

@@ 1 Introduction
Knowledge graphs (KGs) are an integral component of modern information management systems for storing, processing, and managing data. Informally, a KG is a finite collection of facts representing different relations between pairs of nodes, which is typically highly incomplete (Toutanova and Chen, 2015). Motivated by the incompleteness of modern KGs, a new setup for classical query answering has emerged (Ren et al., 2020; Ren and Leskovec, 2020; Zhu et al., 2022; Bai et al., 2023; Yin et al., 2024; Wang et al., 2022; Xu et al., 2023), where the goal is to predict answers that do not necessarily appear in the KG, but are potentially present in its completion. This task is commonly referred to as complex query answering (CQA), and poses a significant challenge, going beyond the capabilities of classical query engines, which typically assume every fact missing from the observable KG is incorrect, following closed-world assumption (Libkin and Sirangelo, 2009).
In its current form, CQA is formulated as a ranking problem: given an input query $Q(x)$ over a KG $G$, the objective is to rank all possible answers based on their likelihood of being a correct answer. Unfortunately, this setup suffers from various limitations. Firstly, this evaluation becomes infeasible for cases where multiple free variables are allowed [^1] As a result, almost all existing proposals focus on queries with only one free variable.. Moreover, to avoid explicitly enumerating solutions, existing methods need to resort to various heuristics and most of them can only handle tree-like queries (Arakelyan et al., 2020; Bai et al., 2023; Zhu et al., 2022) or incur an exponential overhead in more general cases (Yin et al., 2024). Consequently, the structural oversimplification of queries is also reflected in the existing benchmarks. We argue for an alternative problem formulation, more aligned with classical setup, to alleviate these problems.
Problem setup. In this work, we deviate from the existing ranking-based setup, and instead propose and study two query answering problems based on classification. Our first task of interest, query answer classification, involves classifying solutions to queries over knowledge graphs, as $\mathsf{true}$ or $\mathsf{false}$. The second objective, query answer retrieval, requires predicting a correct answer to the query or deciding none exists.
Let us illustrate these tasks on a knowledge graph $G_{\text{ex}}$ (Figure˜1), representing relationships between actors, movies, and locations. The dashed edges denote the missing facts from $G_{\text{ex}}$ and we write $\tilde{G}_{\text{ex}}$ to denote the complete version of $G_{\text{ex}}$ which additionally includes all missing facts. Consider the following first-order query:
@@
Approach and contributions. To solve these tasks, we introduce AnyCQ [^2] The code and data can be found in this GitHub repository., a graph neural network that can predict the satisfiability of a Boolean query over any (incomplete) KG, provided with a function assessing the truth of unobserved links. AnyCQ acts as a search engine exploring the space of assignments to the free and existentially quantified variables in the query, eventually identifying a satisfying assignment to the query. AnyCQ can handle any existentially quantified first-order query in conjunctive or disjunctive normal form. Our contributions can be summarized as follows:
@@ 2 Related work
Link prediction. Earlier models for link prediction (LP) on knowledge graphs, such as TransE (Bordes et al., 2013), RotatE (Sun et al., 2019), ComplEx (Trouillon et al., 2016) and BoxE (Abboud et al., 2020), focused on learning fixed embeddings for seen entities and relations, thus confining themselves to transductive setting. Later, graph neural networks (GNNs) emerged as powerful architectures, with prominent examples including RGCNs (Schlichtkrull et al., 2018) and CompGCNs (Vashishth et al., 2020). These models adapt the message-passing paradigm to multi-relational graphs, thus enabling inductive link prediction on unseen entities. Building on this, Zhu et al. (2021) designed NBFNets which exhibited strong empirical performance via conditional message passing due to its enhanced expressivity (Huang et al., 2023). Recently, ULTRA (Galkin et al., 2024a) became one of the first foundation models on LP over both unseen entities and relations.
@@
Complex query answering. Complex query answering (CQA) (Ren et al., 2020; Ren and Leskovec, 2020) extends the task of link prediction to a broader scope of first-order formulas with one free variable, considering queries with conjunctions ($\land$), disjunctions ($\lor$) and negations ($\neg$). Neuro-symbolic models decompose the CQA task into a series of link prediction problems and employ fuzzy logic to aggregate these individual results. CQD (Arakelyan et al., 2020) was among the first to use fuzzy logic for CQA, relying on pre-trained embeddings and beam search for approximate inference. QTO (Bai et al., 2023) improved on this by exploiting the sparsity of neural score matrices to compute exact solutions without approximation. GNN-QE (Zhu et al., 2022) further advanced the field by training directly over queries, without relying on pre-trained embeddings. FIT (Yin et al., 2024) extended the methodology introduced in QTO to queries containing cycles, at the cost of high complexity. More recently, UltraQuery (Galkin et al., 2024b) has been proposed, leveraging the CQA framework of GNN-QE with ULTRA as the link predictor, achieving the first foundation model for CQA, capable of zero-shot inference on unseen KGs. Neural methods generally rely on neural networks to deduce relations and execute logical connectives simultaneously. CQD-CO (Arakelyan et al., 2020) formulates query answering as a continuous optimization problem: it assigns embeddings to variables with a pre-trained link predictor and optimizes the fuzzy logic score via gradient descent. LMPNN (Wang et al., 2022) employs a novel logical message-passing scheme, leveraging existing KG embeddings to conduct one-hop inferences on atomic formulas. CLMPT (Zhang et al., 2024) extends this methodology by using an attention-based mechanism to aggregate messages incoming from neighbor nodes. Q2T (Xu et al., 2023) utilized the adjacency matrix of the query graph as an attention mask in Transformers (Vaswani et al., 2017) model. While flexible, these methods lack interpretability and variable grounding, and underperform with growing query graph size.
@@
Combinatorial reasoning. GNNs have emerged as a powerful tool for solving combinatorial optimization problems (Cappart et al., 2021). Their power to leverage the inherent structural information encoded in graph representations of instances has been successfully utilized for solving various combinatorial tasks (Joshi et al., 2019; Lemos et al., 2019; Bosnić and Šikić, 2023). As a method of our particular interest, ANYCSP Tönshoff et al. (2023), introduced a novel computational graph representation for arbitrary constraint satisfaction problems (CSP), demonstrating state-of-the-art performance on MAX-CUT, MAX-k-SAT and k-COL.
@@
In this work, we identify answering conjunctive queries as a CSP, tailoring the ANYCSP framework to suit the task of deciding the satisfiability of Boolean formulas over incomplete KGs. Particularly, we integrate link predictors into our architecture to account for the necessity of inferring relations missing in the observable data. We also devise new guidance mechanisms to navigate the search during the early stages, targeting the large domain size. The augmented framework, named AnyCQ, inherits the extrapolation and generalization strength of ANYCSP, resulting in an efficient and effective model for query answer classification and retrieval.
@@ 3 Preliminaries
First-order logic.term is either a constant or a variable. A (binary) atom is an expression of the form $r(t_{1},t_{2})$, where $r$ is a binary relation, and $t_{1},t_{2}$ are terms. A fact, or a ground atom, has only constants as terms. A literal is an atom or its negation. A variable in a formula is quantified (or bound) if it is in the scope of a quantifier; otherwise, it is free. A Boolean formula is a formula without any free variables. A quantifier-free formula is a formula that does not use quantifiers. For notational convenience, we write $\vec{x}=x_{1},...,x_{k}$ and $\vec{y}=y_{1},...,y_{l}$ to represent sequences of variables and $\Phi(\vec{x},\vec{y})$ to represent a quantifier-free formula $\Phi$ using variables from $\{\vec{x},\vec{y}\}$. Similarly, we write $\vec{a}$ to represent tuples of constants of the form $\vec{a}=a_{1},...,a_{k}$. For a first-order logic formula $\Phi(\vec{x})$ with $k$ free variables, we use the notation $\Phi(\vec{a}/\vec{x})$ to represent the Boolean formula obtained by substitution of each free occurrence of $x_{i}$ for $a_{i}$, for all $i$.
Query answering. The focus of this work is on conjunctive queries, i.e., existentially quantified first-order formulas. A conjunctive query (CQ) is a first-order formula of the form $Q(\vec{x})=\exists\vec{y}\,\Phi(\vec{x},\vec{y})$, where $\Phi(\vec{x},\vec{y})$ is a conjunction of literals using variables from $\{\vec{x},\vec{y}\}$. We reserve $\{\vec{y}\}$ for existentially quantified variables and $\{\vec{x}\}$ for free variables. If the query is Boolean, we write $Q=\exists\vec{y}\,\Phi(\vec{y})$.
@@
Given a KG $G$ and a query $Q(\vec{x})=\exists\vec{y}\,\Phi(\vec{x},\vec{y})$, the assignments $\nu:\{\vec{x}\}\to V(G)$, $\mu:\{\vec{y}\}\to V(G)$ respectively map the free and quantified variables to constants. For notational convenience, we denote with $\vec{x}\to{\vec{a}}$ the assignment ${x}_{1}\to{a}_{1}$, …, ${x}_{k}\to{a}_{k}$. We represent by $\Phi(\vec{a}/\vec{x},\vec{e}/\vec{y})$ the formula obtained by substituting the variables with constants according to the assignments $\vec{x}\to{\vec{a}}$ and $\vec{y}\to\vec{e}$. We write $\nu_{x\rightarrow a}$ for an assignment such that $\nu_{x\rightarrow a}(x)=a$ and $\nu_{x\rightarrow a}(z)=\nu(z)$ whenever $z\neq x$.
@@
A Boolean query $Q=\exists\vec{y}\,\Phi(\vec{y})$ evaluates to $\mathsf{true}$ on $G$, denoted $G\models Q$, if there exists an assignment $\vec{y}\to\vec{e}$ such that all positive facts that appear in $\Phi(\vec{e}/\vec{y})$, appear in the set $E(G)$ and none of the negated facts that appear in $\Phi(\vec{e}/\vec{y})$ are present in $E(G)$. In this case, the assignment $\vec{y}\to\vec{e}$ is called a match. For a query $Q(\vec{x})=\exists\vec{y}\,\Phi(\vec{x},\vec{y})$, an assignment $\vec{x}\to\vec{a}$ is called an answer if ${G}\models Q(\vec{a}/\vec{x})$.
In our study, we make a distinction between easy and hard answers to queries, depending on whether the answers can already be obtained from the observed KG $G$ or only from its completion $\tilde{G}$. Formally, given an observed KG $G$ and its completion $\tilde{G}$, we say that an answer $\vec{a}$ is easy (or trivial) if $G\models Q(\vec{a}/\vec{x})$. If, however, $\tilde{G}\models Q(\vec{a}/\vec{x})$ while $G\nvDash Q(\vec{a}/\vec{x})$ then the answer is hard (or non-trivial).
@@
$S_{\pi,G}(r(a,b))$ $=\pi(r,a,b)$
$S_{\pi,G}(\neg Q)$ $=1-S_{\pi,G}(Q)$
$S_{\pi,G}(Q\land Q^{\prime})$ $=\min_{\>}(S_{\pi,G}(Q),S_{\pi,G}(Q^{\prime}))$
$S_{\pi,G}(Q\lor Q^{\prime})$ $=\max_{\>}(S_{\pi,G}(Q),S_{\pi,G}(Q^{\prime}))$
$S_{\pi,G}(\exists x.Q^{\prime}(x))$ $=\max_{a\in V(G)}S_{\pi,G}(Q^{\prime}(a/x))$
@@ 4 Query answering on incomplete KGs
4.1 Limitations of existing problem formulations
Intractability of high-arity query evaluation. The objective of complex query answering is to rank all possible answers to a given logical formula. Already for queries $Q(x_{1},x_{2})$ with two free variables, this entails scoring $|V(G)|^{2}$ pairs of entities $(a_{1},a_{2})\in V(G)^{2}$, which is computationally infeasible for modern knowledge graphs (Toutanova et al., 2015; Carlson et al., 2010) containing thousands of nodes. As a result, most of the existing approaches are not designed to handle higher arity queries, either resolving to inefficient enumeration strategies or approximating answers by marginal predictions. This scalability bottleneck has already been observed by Yin et al. (2023), who suggested more tractable evaluation methodologies, yet again being only marginal approximations of the true performance. Therefore, we argue that the ranking-based formulation has significantly limited the progress in query answering over formulas with multiple free variables.
Limited structural complexity in existing benchmarks. A related limitation lies in the structural simplicity of existing benchmarks (Ren and Leskovec, 2020). Standard CQA literature predominantly focuses on tree-like queries, which aligns with the capabilities of most current models. More recently, Yin et al. (2023) introduced a dataset containing cyclic queries and queries with up to two free variables; however, the overall structures remained constrained - featuring at most four variables and a single cycle. We argue that addressing structurally richer queries is essential for advancing automated reasoning systems. In real-world applications to autonomous systems, such as an AI trip planner that simultaneously books flights, accommodations, and activities while satisfying budget and availability constraints, the underlying reasoning involves large, complex queries with multiple variables. As AI agents become more capable, the complexity of the queries they must resolve is only expected to increase, hence requiring more expressive answering engines.
Lack of probabilistic calibration in ranking-based methods. Practical applications often demand binary decisions - answering questions like “Is X true?” or “What is the correct answer to Y?”, requiring models to classify candidate solutions as either true or false (van Bakel et al., 2021). However, many ranking-based CQA methods do not natively support this decision-making paradigm, as they focus on ordering candidates without enforcing a meaningful threshold to distinguish valid answers from incorrect ones. While many of these models are trained using classification losses, such training does not guarantee the output scores correspond to calibrated satisfiability probabilities. In fact, several approaches rely on Noisy Contrastive Estimation (Wang et al., 2022; Zhang et al., 2024) or apply ad hoc score-to-probability transformations (Bai et al., 2023; Yin et al., 2024), further weakening the reliability of predicted scores in downstream tasks.
@@ 4.2 Query Answer Classification & Query Answer Retrieval
To address these limitations, we propose two new query answering tasks designed to provide more targeted responses while ensuring scalability for more complex logical queries.
@@
Query answer retrieval assesses the correctness of the top-ranked result. By requiring models to either deliver a correct assignment to the free variables of the input query or confidently assert the absence of one, QAR aligns more closely with practical decision-making processes, ensuring the output is both relevant and reliable:
@@
Reduction to conjunctive query answering. Following earlier works (Wang et al., 2022; Yin et al., 2024; Zhang et al., 2024), we note that each existential first-order logic (EFO) query $Q(\vec{x})$ can be represented as an equivalent $Q^{\prime}(\vec{x})\equiv Q(\vec{x})$ in disjunctive normal form: $Q^{\prime}(\vec{x})=\exists\vec{y}.(D_{1}(\vec{x},\vec{y})\lor...\lor D_{n}(\vec{x},\vec{y}))$, where each $D_{i}$ is a conjunction of literals. Under this decomposition, $Q^{\prime}(\vec{x})\equiv Q_{1}(\vec{x})\lor...\lor Q_{n}(\vec{x})$, where $Q_{i}(\vec{x})=\exists\vec{y}.D_{i}(\vec{x},\vec{y})$. Hence $Q(\vec{a}/\vec{x})$ is satisfiable if and only if one of $Q_{i}(\vec{a}/\vec{x})$ is satisfiable, reducing the problem to conjunctive queries. Contrary to the ranking setting, which requires combining the results for components $Q_{i}(\vec{a}/\vec{x})$ to obtain a score for $Q(\vec{a}/\vec{x})$, in the classification-based formulation, this aggregation considers only binary ‘true’ and ‘false’ labels. Consequently, we concentrate on conjunctive queries, since answering $Q(\vec{x})$ is equivalent to solving each $Q_{i}(\vec{x})$ individually, and the need for merging the outcomes does not introduce further complexity.

### v3

@@ 1 Introduction
Knowledge graphs (KGs) are an integral component of modern information management systems for storing, processing, and managing data. Informally, a KG is a finite collection of facts representing different relations between pairs of nodes. Motivated by the incompleteness of modern KGs [toutanova2015observed], a new setup for classical query answering has emerged [ren2020query2box, betae, gnn-qe, qto, fit, lmpnn, q2t], where the goal is to predict answers that do not necessarily appear in the KG, but are potentially present in its completion. This task is commonly referred to as complex query answering (CQA), and poses a significant challenge, going beyond the capabilities of classical query engines, which typically assume every fact missing from the observable KG is incorrect, following closed-world assumption [open_world_assumption].
In its current form, CQA is formulated as a ranking problem: given an input query $Q(x)$ over a KG $G$, the objective is to rank all possible answers based on their likelihood of being a correct answer. Unfortunately, this setup suffers from various limitations. Firstly, this evaluation becomes infeasible for cases where multiple free variables are allowed [^1] As a result, almost all existing proposals focus on queries with only one free variable.. Moreover, to avoid explicitly enumerating solutions, existing methods need to resort to various heuristics and most of them can only handle tree-like queries [cqd, qto, gnn-qe] or incur an exponential overhead in more general cases [fit]. Consequently, the structural oversimplification of queries is also reflected in the existing benchmarks. We argue for an alternative problem formulation, more aligned with classical setup, to alleviate these problems.
Problem setup. In this work, we deviate from the existing ranking-based setup, and instead propose and study two query answering problems based on classification. Our first task of interest, query answer classification, involves classifying solutions to queries over knowledge graphs, as $\mathsf{true}$ or $\mathsf{false}$. The second objective, query answer retrieval, requires predicting a correct answer to the query or deciding that none exists.
Example. Let us illustrate these tasks on a knowledge graph $G_{\text{ex}}$ (Figure˜1), representing relationships between actors, movies, and locations. The dashed edges denote the missing facts from $G_{\text{ex}}$ and we write $\tilde{G}_{\text{ex}}$ to denote the complete version of $G_{\text{ex}}$ which additionally includes all missing facts. Consider the following first-order query:
@@
Approach and contributions. To solve these tasks, we introduce AnyCQ, a graph neural network that provided with a function assessing the truth of unobserved links, can predict the satisfiability of a Boolean query over any (incomplete) KG. AnyCQ acts as a search engine exploring the space of assignments to the free and existentially quantified variables in the query, eventually identifying a satisfying assignment to the query. AnyCQ can handle any existentially quantified first-order query in conjunctive or disjunctive normal form. Our contributions can be summarized as follows:
@@ 2 Related work
Link prediction. Earlier models for link prediction (LP) on knowledge graphs, such as TransE [transe], RotatE [sun2019rotate], ComplEx [trouillon2016complex], and BoxE [abboud2020boxe], learn fixed embeddings for entities and relations, confining themselves to transductive setting. Later, graph neural networks (GNNs) emerged as powerful architectures, with prominent examples including RGCNs [schlichtkrull2018modeling] and CompGCNs [vashishth2020composition]. These models adapt the message-passing paradigm to multi-relational graphs, thus enabling inductive link prediction on unseen entities. Building on this, nbfnet proposed NBFNets, which achieve strong performance through conditional message passing [huang2023a]. Recently, ULTRA [galkin2023ultra] became one of the first foundation models on LP over both unseen entities and relations.
@@
Complex query answering. Complex query answering (CQA) [ren2020query2box, betae] generalizes link prediction to first-order formulas with one free variable, considering queries with conjunctions ($\land$), disjunctions ($\lor$) and negations ($\neg$). Neuro-symbolic models decompose the CQA task into a series of link prediction problems, combining results with fuzzy logic. CQD [cqd] pioneered this approach with beam search over pre-trained embeddings for approximate inference. QTO [qto] improved on this by exploiting the sparsity of neural score matrices to compute exact solutions without approximation. FIT [fit] extended QTO to cyclic queries at a higher cost. GNN-QE [gnn-qe] trained directly over queries without relying on pre-trained embeddings. UltraQuery [galkin2024ultraquery] combined GNN-QE’s framework with ULTRA, yielding the first foundation model for CQA with zero-shot generalization. Neural methods generally rely on neural networks to deduce relations and execute logical connectives simultaneously. CQD-CO [cqd] formulates query answering as continuous optimization, assigning embeddings to variables and optimizing the fuzzy logic with gradient descent. LMPNN [lmpnn] and CLMPT [clmpt] employ logical message-passing and attention-based aggregation. Q2T [q2t] utilized the adjacency matrix of the query graph as an attention mask in Transformers [attention_is_all] model. While flexible, these methods lack interpretability and variable grounding, and underperform with growing query graph size.
@@
Combinatorial reasoning. GNNs have emerged as a powerful tool for solving combinatorial optimization problems [gnn_co_survey]. Their power to leverage the inherent structural information encoded in graph representations of instances has been successfully utilized for solving various combinatorial tasks [joshi, decision_col, hamiltonian_cycles]. As a method of our particular interest, ANYCSP [anycsp], introduced a novel computational graph representation for arbitrary constraint satisfaction problems (CSP), demonstrating state-of-the-art performance on MAX-CUT, MAX-k-SAT and k-COL.
@@
In this work, we cast conjunctive queries as a CSP, tailoring the ANYCSP framework to suit the task of satisfiability of Boolean formulas over incomplete KGs. AnyCQ integrates link predictors to infer missing relations and introduces guidance mechanisms for efficient search over large domains. Leveraging ANYCSP’s extrapolation and generalization strengths, AnyCQ provides an effective solution for query answer classification and retrieval.
@@ 3 Preliminaries
First-order logic.term is either a constant or a variable. A (binary) atom is an expression of the form $r(t_{1},t_{2})$, where $r$ is a binary relation, and $t_{1},t_{2}$ are terms. A fact, or a ground atom, has only constants as terms. A literal is an atom or its negation. A variable in a formula is quantified (or bound) if it is in the scope of a quantifier; otherwise, it is free. A Boolean formula has no free variables. A quantifier-free formula does not use quantifiers. We write $\vec{x}=x_{1},...,x_{k}$ and $\vec{y}=y_{1},...,y_{l}$ to represent sequences of variables and $\Phi(\vec{x},\vec{y})$ to represent a quantifier-free formula $\Phi$ using variables from $\{\vec{x},\vec{y}\}$. Similarly, we write $\vec{a}$ to represent tuples of constants of the form $\vec{a}=a_{1},...,a_{k}$. For a first-order logic formula $\Phi(\vec{x})$ with $k$ free variables, we use $\Phi(\vec{a}/\vec{x})$ to represent the Boolean formula obtained by substitution of each free occurrence of $x_{i}$ for $a_{i}$, for all $i$.
@@
Query answering. A conjunctive query (CQ) is a first-order formula of the form $Q(\vec{x})=\exists\vec{y}\,\Phi(\vec{x},\vec{y})$, where $\Phi(\vec{x},\vec{y})$ is a conjunction of literals using variables from $\{\vec{x},\vec{y}\}$. We reserve $\{\vec{y}\}$ for existentially quantified variables and $\{\vec{x}\}$ for free variables. If the query is Boolean, we write $Q=\exists\vec{y}\,\Phi(\vec{y})$. Given a KG $G$ and a query $Q(\vec{x})=\exists\vec{y}\,\Phi(\vec{x},\vec{y})$, the assignments $\nu:\{\vec{x}\}\to V(G)$, $\mu:\{\vec{y}\}\to V(G)$ respectively map the free and quantified variables to constants. For notational convenience, we denote with $\vec{x}\to{\vec{a}}$ the assignment ${x}_{1}\to{a}_{1}$, …, ${x}_{k}\to{a}_{k}$. We represent by $\Phi(\vec{a}/\vec{x},\vec{e}/\vec{y})$ the formula obtained by substituting the variables with constants according to the assignments $\vec{x}\to{\vec{a}}$ and $\vec{y}\to\vec{e}$. We write $\nu_{x\rightarrow a}$ for an assignment such that $\nu_{x\rightarrow a}(x)=a$ and $\nu_{x\rightarrow a}(z)=\nu(z)$ whenever $z\neq x$.
@@
A Boolean query $Q=\exists\vec{y}\,\Phi(\vec{y})$ evaluates to $\mathsf{true}$ on $G$, denoted $G\models Q$, if there exists an assignment $\vec{y}\to\vec{e}$ such that all positive facts that appear in $\Phi(\vec{e}/\vec{y})$, appear in the set $E(G)$ and none of the negated facts that appear in $\Phi(\vec{e}/\vec{y})$ are present in $E(G)$. In this case, the assignment $\vec{y}\to\vec{e}$ is called a match. For a query $Q(\vec{x})=\exists\vec{y}\,\Phi(\vec{x},\vec{y})$, an assignment $\vec{x}\to\vec{a}$ is called an answer if ${G}\models Q(\vec{a}/\vec{x})$. We distinguish between easy and hard answers. An answer $\vec{a}$ is easy (or trivial) if $G\models Q(\vec{a}/\vec{x})$; it is hard (or non-trivial) if $\tilde{G}\models Q(\vec{a}/\vec{x})$ but $G\nvDash Q(\vec{a}/\vec{x})$.
@@
[…]
@@ 4 Query answering on incomplete KGs
Existing problem formulations for complex query answering (CQA) suffer from several fundamental limitations that restrict directions for progress. First, the standard ranking-based objective is computationally infeasible for queries with multiple free variables, as it requires scoring all candidate answers. Their number is exponential in the number of free variables, which is a known bottleneck for large knowledge graphs [fb15k237, nell, efok_dataset]. This has also led to reliance on benchmarks with limited structural complexity, featuring either tree-like queries [betae] or ones with a single cycle [fit], failing to capture the richness of real-world reasoning tasks. Finally, many ranking-based models are not probabilistically calibrated for binary classification [cqa_classification]. Their scores often result from non-probabilistic training objectives [lmpnn, clmpt] or require ad-hoc transformations [qto, fit], making them unsuitable for applications that demand definitive true/false decisions. See extended discussion in Section˜A.1.
@@ 4.1 Query Answer Classification & Query Answer Retrieval
In this section, we propose two new query answering tasks designed to provide more targeted responses to complex logical queries, while avoiding the complexity incurred by answer enumeration.
@@
Query answer retrieval assesses the correctness of the top-ranked result. By requiring models to either deliver a correct assignment to the free variables of the input query or assert the absence of one, QAR aligns more closely with practical decision-making, ensuring the output is relevant and reliable:
@@
Reduction to conjunctive query answering. As shown in prior work [lmpnn, fit, clmpt], any existentially quantified first-order (EFO) query $Q(\vec{x})$ can be rewritten in disjunctive normal form $Q^{\prime}(\vec{x})=\exists\vec{y}\,(D_{1}(\vec{x},\vec{y})\lor\cdots\lor D_{n}(\vec{x},\vec{y})),$ where each $D_{i}$ is a conjunction of literals. Thus, we have $Q^{\prime}(\vec{x})\equiv Q_{1}(\vec{x})\lor\cdots\lor Q_{n}(\vec{x})$ with $Q_{i}(\vec{x})=\exists\vec{y}\,D_{i}(\vec{x},\vec{y})$, and $Q(\vec{a}/\vec{x})$ is satisfiable if and only if some $Q_{i}(\vec{a}/\vec{x})$ is satisfiable, reducing the task to conjunctive queries. Unlike ranking, which must combine scores across disjuncts, classification aggregates only binary outcomes. Hence we focus on conjunctive queries, as solving each $Q_{i}$ individually introduces no additional complexity.
5 AnyCQ: framework for query answering
To address the introduced tasks of query answer classification and retrieval, we propose a neuro-symbolic framework for scoring arbitrary existential Boolean formulas, called AnyCQ. Let $\pi$ be a link predictor for an observable knowledge graph $G$. An AnyCQ model $\Theta$ equipped with $\pi$ can be viewed as a function $\Theta(G,\pi):\mathsf{CQ^{0}}(G)\rightarrow[0,1]$ where $\mathsf{CQ^{0}}(G)$ is the class of conjunctive Boolean queries over the same vocabulary as $G$. For input $Q=\exists\vec{y}.\Phi(\vec{y})$, $\Theta$ searches over the space of assignments to $\vec{y}$ for
and returns an approximation $\Theta(Q|G,\pi)$ of $S_{\pi,G}(Q)$ as
Note that by unfolding the Boolean formula score: $S_{\pi,G}(Q)=S_{\pi,G}(\Phi(\alpha_{\max}(\vec{y})/\vec{y}))\approx\Theta(Q|G,\pi).$ Hence, by leveraging the potential of GNNs for solving combinatorial optimization problems, we can recover strong candidates for $\alpha_{\max}$, allowing for an accurate estimation of $S_{\pi,G}(Q)$.
Overview. During the search, our method encodes the query $Q$ and its relation to the current assignment $\alpha$ into a computational graph $G_{Q,\alpha}$ (Section˜5.1). This graph is then processed with a simple GNN $\theta$ (whose architecture is described in Section˜B.1), which updates its hidden embeddings and generates distributions $\mu$ from which the next assignment $\alpha^{\prime}$ is sampled (Section˜5.2).
5.1 Query representation
We transform the input queries into a computational graph (Figure˜2), whose structure is adopted from ANYCSP [anycsp]. Consider a conjunctive Boolean query $Q=\exists\vec{y}.\Phi(\vec{y})$ over a knowledge graph $G$, with $\Phi$ quantifier-free, and let $\pi$ be a link predictor for $G$. Let $c_{1},...,c_{n}$ be constant symbols mentioned in $\Phi$, and $\psi_{1},...,\psi_{l}$ be the literals in $\Phi$. We define the domain ${\mathcal{D}}(e)$ of the term $e$ as ${\mathcal{D}}(y)=V(G)$ for each existentially quantified variable $y$ and ${\mathcal{D}}(c_{i})=\{c_{i}\}$ for each constant $c_{i}$. Given an assignment $\alpha:\vec{y}\rightarrow\vec{a}$, the computational graph $G_{Q,\alpha}$ is constructed as follows: